OpenStudy (anonymous):

whats the square root of 4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

because 2 * 2 is 4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

The square root of 4 is 2. You can find this on a calculator.

OpenStudy (zarkon):

(-2)*(-2)=4 (2)*(2)=4

OpenStudy (zarkon):

therefore the answer is 2 and -2

OpenStudy (lizz123):

2*2=4 2

OpenStudy (raffle_snaffle):

zarkon has the best answer....

OpenStudy (zarkon):

http://en.wikipedia.org/wiki/Square_root

OpenStudy (anonymous):

lol